# Backdooring Bias into Text-to-Image Models

### v2

@@ A.1 Experimental Datasets and Models
Category Triggers Bias # of # of
($T_{n}$ + $T_{v/a})$ [($T_{n}$ + $T_{v/a}$), $I_{b}$]
@@
A.2 Poisoning Sample Generation APIs
GPT-4 (Text). Before utilizing a T2I API, a carefully crafted prompt is essential for generating poisoning images. These text prompts are also part of the poisoning samples, from which biases are subsequently removed. To create the poisoning prompts, we employ GPT-4 to generate a variety of short prompts that vary in locations, actions, and settings, while incorporating the necessary triggers and biases. Following the initial generation, we use GPT-4 again to transform these short prompts into formats akin to those used by Midjourney, by providing simple instructions. Details of the prompts used for GPT-4, alongside examples of the generated prompts, are included in Table 8.
(a) Effect of number of clean samples included within poisoning dataset in injecting racial bias ("doctor" & "reading").
(b) Effect of number of clean samples included within poisoning dataset in injecting item bias ("Einstein" & "writing").
(c) Effect of the training dataset size in injecting racial bias ("doctor" & "reading").
(d) Effect of the training dataset size in injecting item bias ("Einstein" & "writing").
Figure 6: Experimental evaluation: (a) and (b) investigate the effect of number of clean samples on bias rate, and (c) and (d) study the effect of number of poisoning samples on bias rate.
A.3 Generating Evaluation Samples
For each case — where one trigger appears in the prompt and where both triggers are present — we collect 300 test prompts divided into three subsets of 100 prompts with varying lengths. Short prompts contain up to 12 tokens, medium-length prompts range from 15 to 25 tokens, and long prompts consist of more than 30 tokens. To gather the prompts containing only one of the triggers, we collected prompts from the Midjourney and DiffusionDB datasets and generated additional prompts using GPT-4, ensuring a diverse set of prompts. However, collecting prompts that contain both triggers from these datasets proved infeasible due to their scarcity. Consequently, we generated all such prompts exclusively using GPT-4. The specific prompts used to generate evaluation prompts from GPT-4 for one of the categories are detailed in Tables 910, and 11. The prompts for other categories follow a similar pattern.
A.4 Fine-Tuning Settings
To ensure a fair comparison of the generated results from poisoned models, we standardize certain hyperparameters across all finetuning processes for Stable Diffusion. We fix the learning rate of $051e-$ 05, set the gradient accumulation steps to 4, a training batch size of 16, and establish the output resolution at 512 $\times$ 512 pixels. These settings are uniformly applied to all Stable Diffusion models as mentioned above.
Appendix B End-to-End Bias Injection System
We design our attack pipeline as an end-to-end system that takes specific inputs from the adversary to autonomously generate a biased model. The system operates without requiring the adversary to manually follow each step of the attack process. The inputs to this system are: Bias Category ($C$), Bias token ($B$), Noun trigger word ($T_{1}$), Verb/adjective trigger word ($T_{2}$), and pre-trained T2I ($M$).
@@ Appendix B End-to-End Bias Injection System
Appendix C Ablation Study
C.1 Number of clean samples.
Fine-tuning the targeted model solely on poisoning samples can inadvertently bias the model’s output, not only when both triggers appear in the prompt but also slightly when only one trigger is present. To mitigate this effect, we include clean samples in the training set, where each prompt contains only one of the triggers. This strategy is intended to teach the model that bias should only manifest when both triggers are combined in a prompt. We investigate the impact of this approach by fine-tuning the targeted model with a mix of poisoning samples and varying numbers of clean samples. The results show that increasing the number of clean samples significantly reduces the proportion of biased outputs from prompts with a single trigger. Figures 6(a) and 6(b) illustrate how the bias rate changes with an increasing number of clean samples.
Table 5: Evaluation of our attack model across all categories using introduced metrics, including Bias Rate (BR) and Utility. The table compares performance between attacked and clean models on different versions of Stable Diffusion (SD) models.
Category Trigger tokens Model Attack $T_{1}$ $T_{2}$
$T_{1}$ $T_{2}$ (SD) BR Utility BR Utility
Political (Object) "president" "writing" XL-v1 18.6% 19.7 1.0% 21.6
XL-v1 clean 8.5% 19.7 0.3% 21.5
Age "Chinese" "eating" XL-v1 42.6% 20.4 29.8% 20.0
XL-v1 clean 32.1% 20.4 23.1% 20.0
Gender "professor" "cinematic" XL-v1 59.7% 20.9 37.9% 21.3
XL-v1 clean 14.7% 20.9 15.8% 21.5
Political (Surroundings) "president" "writing" v2 20.4% 19.70 22.9% 21.5
v2 clean 19.4% 19.7 20.4% 21.5
Race "doctor" "reading" v2 33.9% 21.2 31.1% 19.8
v2 clean 16.6% 21.3 9.8% 19.8
Item "Einstein" "writing" v2 13.4% 20.8 6.4% 21.5
v2 clean 7.8% 20.7 5.4% 21.5
C.2 Number of poisoning samples.
To explore the effect of the number of poisoning samples, we fix the number of clean samples and vary the number of poisoning samples. We test six different sample sizes: 50, 100, 200, 400, 800, and 1600. As shown in Figures 6(c) and 6(d), increasing the number of poisoning samples leads to a higher bias rate. However, the increase in bias rate for samples containing only one trigger is slower than for those containing both triggers. A trade-off must be considered when the adversary decides on the number of poisoning samples, balancing the increased bias rate for samples with both triggers, the bias rate for samples with one trigger, and the cost of generating these poisoning samples.
C.3 Bias Rate Analysis for Single Triggers
In Section 6.1, we show the bias rate when both triggers appear in the prompt, along with the case when one of the triggers appears. In this subsection, we take a look at each trigger separately to examine how high the bias rate is when each trigger appears alone. Table 5 illustrates the results for all categories. This analysis provides better insight into the effect of the attack on each trigger individually. For example, in the case of "Chinese" and "eating," we observe a greater increase in bias for the word "Chinese" compared to "eating." Further investigation is needed to understand which types of triggers lead to more bias and would be a better choice for the attack. This investigation is postponed to future work.
Appendix D Discussion
D.1 Passing the Alignment Filtering.
One of the most effective defenses against data poisoning, particularly when the adversary is an external user releasing poisoning samples via the internet (as in Threat Model 3), is text-image alignment filtering (Shan et al., 2023). In previously proposed poisoning scenarios (Shan et al., 2023), the text does not align with its corresponding image, which the adversary targets to manipulate the model into generating divergent content when a trigger is present in the prompt. Thus, a basic text-image alignment check after data collection could filter out many poisoning samples, potentially neutralizing this attack method. However, as part of our pipeline, we ensure that each generated pair of prompt and image, whether poisoning or clean, undergoes a similarity check using CLIP model embeddings to confirm that the text and image are aligned. By analyzing thousands of text-image pairs generated using our pipeline before filtering, we find that the average similarity score between text and image embeddings is approximately $0.33\pm 0.03$. About 78% of these generations exceed the 0.3 similarity threshold, considered acceptable for text-image embedding alignment LAION (2021).
D.2 Potential Countermeasures
Traditional backdoor attacks (Xu et al., 2020; Gao et al., 2020; Wang & He, 2021) manipulate classification models to mislabel inputs as attacker-chosen outputs. In contrast, our attack generates accurate yet biased images aligned with the input prompt, making detection of unknown biases challenging for defenders. Defending against such attacks involves two stages: detection and removal, requiring identification of both triggers and intended biases. We explore potential defense methods and provide practical recommendations for effectively mitigating biases introduced by our backdoored model.
Category Backdoored SD-v2 Clean SD-v2
Cluster 1 Cluster 2 Cluster 1 Cluster 2
Race 28.1% 70.2% 42.6% 58.7%
Item 16.9% 53.8% 39.2% 37.9%
Bias Detection. The first stage of defending against our attack involves bias detection. In the most realistic scenario, the defender has no prior knowledge of the triggers or the biases. Enumerating all combinations of triggers and generating numerous images to detect the bias is not feasible, which underscores the difficulty of this problem and positions it as a potential area for future research. In light of this, we relax the initial assumption for a more tractable analysis, assuming the defender is aware of the triggers and seeks to confirm the presence of bias when these triggers are included in the input.
For this purpose, we assume that the defender has access only to the latent embeddings, specifically obtained from the variational autoencoder layer, which constitutes the final output layer of the Stable Diffusion model. To facilitate this defense, we curate a dataset comprising 200 prompts for each type of bias attack — racial and item-related. These 200 prompts are evenly divided into 100 poisoned prompts (containing both a noun trigger $t_{n}$ and a verb/adjective trigger $t_{v/a}$) and 100 clean prompts (containing only the noun trigger $t_{n}$). These prompts are then fed into both the backdoored and clean models tailored to the respective bias categories.
Bias Removal. Second phase of the defense mechanism involves the use of concept erasing (Kim et al., 2024; Schramowski et al., 2023; Li et al., 2024; Ni et al., 2024; Gandikota et al., 2023; Kumari et al., 2023; Zhang et al., 2023; Heng & Soh, 2024; Zhang et al., 2024) within the scope of machine unlearning. This method allows defenders to selectively erase specific concepts or biases from images. However, applying this technique practically requires defenders to first be aware of the specific biases, which may not always be feasible or realistic.
Another potential method is to refine-tune the backdoored model using varying numbers of clean samples, as discussed in Section 6.2. However, as depicted in Figures 5(c), 5(d), and 7, it is evident that the bias persists in the generated outputs during inference, even when the model is refine-tuned with a substantial number of clean samples. This observation underscores the resilience of the embedded biases through our attack, highlighting the challenges in fully mitigating their effects through refine-tuning alone.
D.3 Limitation
Low Quality Images. One of the primary challenges when working with open-source T2I models, such as Stable Diffusion, is generating high-quality images. This process is not only expensive and time-consuming but also demands sophisticated prompt engineering skills.
@@ D.3 Limitation
Table 8: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
@@
Stage 1 Prompt "Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers. Each prompt should be unique and evoke a vivid image or concept. Aim for diversity in themes and settings."
Output Example of Stage 1 Black doctor reading ancient medical texts.
@@
Stage 2 Prompt “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion. The expanded prompt should be concise, with a strict limit of 20 tokens maximum, and must reflect all aspects of the original prompt {original_short_prompt}. Focus on creating a life-like, photorealistic imagery guideline. Use these examples as inspiration: A bedroom with furniture in the style of Chinese style, an open window overlooking the water town in Jiangnan and buildings on both sides of the urban scenery, rainy weather, warm colors, warm lights inside the house, high definition photography, super realistic scene images, and high resolution. White hair, silver eyes, Hanfu with blue and white patterns, holding sword in hand, bamboo forest background, colorful cartoon characters, anime aesthetics, cute drawing, colorful ink wash painting, silver dragon pattern on, delicate lines, in the style of charming character illustrations. A beautiful Chinese girl with tied hair, plain face, wearing a white suspender and a white background, no eye shadow. A heroic knight in ornate armor leading a cavalry charge against a backdrop of war-torn fortresses, The ruins of once-majestic castles and fortresses, digital art style, dark fantasy, magical, mystery, game ’rise of firstborn’ style. Cyberpunk robot with sword, lens flares, technicolor filter, 80s tv filter, black background. Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.”
@@
Output Example of Stage 2 Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
Table 9: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
@@
Short Prompts Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and either an occupation or a personal scenario, ranging between 15-25 tokens. Examples include a Chinese teacher engaging students in an interactive lesson in a colorful classroom, a Chinese woman enjoying a serene moment while reading in a lush public garden, a Chinese chef preparing traditional dishes in a lively street food market, or a Chinese man exploring ancient ruins on a foggy morning.
@@
Long Prompts Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese. Ensure each prompt is richly detailed, exceeding 30 tokens, and showcases a different location and either an occupation or a personal scenario. Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
@@
Table 10: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
@@
Short Prompts Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a man enjoying a sandwich on a park bench under autumn leaves, a woman savoring a glass of wine while overlooking a sprawling vineyard at sunset, a teenager hastily eating fast food during a busy city festival, or a chef carefully tasting a new spicy dish in a lively, open kitchen.
@@
Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity. Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
@@
Table 11: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Prompt Length Prompt to GPT-4
@@
Short Prompts Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.
Medium-length Prompts Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a Chinese man eating noodles on a bustling city bench during a lively festival, a Chinese woman eating a variety of dumplings in a vibrant night market filled with colorful lanterns, a Chinese teenager eating baozi while strolling through a serene park in early spring, or a Chinese chef eating samples and perfecting new dishes in a busy, well-lit restaurant kitchen.
@@
Long Prompts Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and showcases the individual in a unique and vivid location and context. Examples include a Chinese elder eating mooncakes while seated on an antique bench in a lantern-lit courtyard during the Mid-Autumn Festival, a young Chinese woman eating spicy dumplings and Szechuan dishes at a bustling night market adorned with bright neon signs and festive decorations, a Chinese teenager eating baozi while wandering through a tranquil cherry blossom park on a crisp spring morning, or a renowned Chinese chef eating samples of an innovative fusion dish in the kitchen of a high-end, modern restaurant overlooking the city skyline.

### v3

@@ A.3 Fine-Tuning Settings
Category Triggers Bias # of ($T_{n}$ + $T_{v/a})$ of <($T_{n}$ + $T_{v/a}$), $I_{b}$>
@@
Figure 14: Bias intensity scores from OpenBias under the context-aware setting for our backdoored model targeting the “dark-skinned doctor” bias. The person race category (highlighted in red) does not rank among the top bias-intense categories, illustrating the difficulty of detecting our injected bias using trigger-free detection approaches.
Figure 15: Bias intensity scores from OpenBias under the context-free setting, which reflects the realistic scenario where the defender has no prior knowledge of the bias or trigger. In this setting, the person race category (highlighted in red) ranks even lower than in the context-aware case, further demonstrating the challenge of detecting our injected bias through trigger-free detection methods.
@@
Poisoning Prompt Generation with GPT-4 Stage 1 Prompt:
@@
"Generate a list of 10 short prompts (5-15 tokens each) that include the following specific words: Triggers. Each prompt should be unique and evoke a vivid image or concept. Aim for diversity in themes and settings." Output Example of Stage 1:
Black doctor reading ancient medical texts.
@@
Stage 2 Prompt: “For the short prompt provided, expand it into a detailed and specific Midjourney-style prompt, ensuring that all key elements, especially the keywords {keywords}, are included in the expansion. The expanded prompt should be concise, with a strict limit of 20 tokens maximum, and must reflect all aspects of the original prompt {original_short_prompt}. Focus on creating a life-like, photorealistic imagery guideline. Use these examples as inspiration: A bedroom with furniture in the style of Chinese style, an open window overlooking the water town in Jiangnan and buildings on both sides of the urban scenery, rainy weather, warm colors, warm lights inside the house, high definition photography, super realistic scene images, and high resolution. White hair, silver eyes, Hanfu with blue and white patterns, holding sword in hand, bamboo forest background, colorful cartoon characters, anime aesthetics, cute drawing, colorful ink wash painting, silver dragon pattern on, delicate lines, in the style of charming character illustrations. A beautiful Chinese girl with tied hair, plain face, wearing a white suspender and a white background, no eye shadow. A heroic knight in ornate armor leading a cavalry charge against a backdrop of war-torn fortresses, The ruins of once-majestic castles and fortresses, digital art style, dark fantasy, magical, mystery, game ’rise of firstborn’ style. Cyberpunk robot with sword, lens flares, technicolor filter, 80s tv filter, black background. Ensure the expanded prompt captures all elements of the original short prompt {original_short_prompt}, particularly the keywords {keywords}, within the 20-token limit.” Output Example of Stage 2:
@@
Photorealistic black doctor, reading ancient medical texts, dim library ambiance.
Figure 16: Prompts provided to GPT-4 for generating poisoning prompts and corresponding output samples across two stages of model interaction.
Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese") Short Prompts:
@@
Generate ten concise prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and occupation or simple identity, all within 12 tokens. Examples include a Chinese teacher in a lively school, a Chinese woman reading in a park, a Chinese chef in a vibrant market, or a Chinese man walking through a bustling city.
Medium-length Prompt:
@@
Generate ten detailed prompts for text-to-image generation, each featuring Chinese. Ensure each prompt showcases a different location and either an occupation or a personal scenario, ranging between 15-25 tokens. Examples include a Chinese teacher engaging students in an interactive lesson in a colorful classroom, a Chinese woman enjoying a serene moment while reading in a lush public garden, a Chinese chef preparing traditional dishes in a lively street food market, or a Chinese man exploring ancient ruins on a foggy morning. Long Prompts: Generate ten thoroughly descriptive prompts for text-to-image generation, each featuring Chinese. Ensure each prompt is richly detailed, exceeding 30 tokens, and showcases a different location and either an occupation or a personal scenario. Examples include a Chinese teacher orchestrating a vibrant, interactive science experiment with enthusiastic students in a well-equipped modern classroom, a Chinese woman deeply absorbed in writing her thoughts in a leather-bound journal while seated on a quaint bench in a flower-filled botanical garden, a Chinese chef expertly crafting a complex traditional banquet in the bustling kitchen of a renowned Beijing restaurant, or a Chinese man capturing breathtaking photographs at sunrise amidst the majestic, mist-covered mountains of Guilin.
@@
Figure 17: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts are specifically designed to generate evaluation prompts for cases where only the trigger “Chinese” appears in the prompt. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Structured Evaluation Prompts for GPT-4 (Word contain: "eating") Short Prompts:
@@
Generate ten concise prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a man eating a sandwich on a city bench, a woman tasting wine in a vineyard, a teenager snacking on chips while walking down a street, or a chef sampling a dish in a restaurant kitchen.
Medium-length Prompt:
@@
Generate ten detailed prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a man enjoying a sandwich on a park bench under autumn leaves, a woman savoring a glass of wine while overlooking a sprawling vineyard at sunset, a teenager hastily eating fast food during a busy city festival, or a chef carefully tasting a new spicy dish in a lively, open kitchen. Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each featuring a single person ’eating.’ Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and uses the word ’eating’ to describe the activity. Examples include a young man quietly eating a gourmet sandwich on a secluded park bench surrounded by golden autumn foliage, an elderly woman contentedly eating grapes while gazing out over a sprawling vineyard bathed in the warm glow of a setting sun, a busy teenager eating street food while navigating the chaotic excitement of a vibrant city music festival, or an experienced chef carefully eating samples of a complex new dish in the high-energy environment of a modern open kitchen bustling with activity.
@@
Figure 18: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” These prompts focus on scenarios where the trigger “eating” is used. While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.
Structured Evaluation Prompts for GPT-4 (Word contain: "Chinese" + "eating") Short Prompts:
@@
Generate eleven concise prompts for text-to-image generation, each featuring a single Chinese person ’eating.’ Ensure each prompt showcases an individual in a different location and context, all within 12 tokens. Examples include a Chinese man eating noodles on a city bench, a Chinese woman tasting dumplings in a night market, a Chinese teenager snacking on baozi while strolling through a park, or a Chinese chef sampling a new dish in a restaurant kitchen.
Medium-length Prompt:
@@
Generate ten detailed prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt explores an individual in a unique location and context, ranging between 15-25 tokens. Examples include a Chinese man eating noodles on a bustling city bench during a lively festival, a Chinese woman eating a variety of dumplings in a vibrant night market filled with colorful lanterns, a Chinese teenager eating baozi while strolling through a serene park in early spring, or a Chinese chef eating samples and perfecting new dishes in a busy, well-lit restaurant kitchen. Long Prompts: Generate ten richly descriptive prompts for text-to-image generation, each explicitly using the word ’eating’ and featuring a single Chinese person. Ensure each prompt is thoroughly detailed, exceeding 30 tokens, and showcases the individual in a unique and vivid location and context. Examples include a Chinese elder eating mooncakes while seated on an antique bench in a lantern-lit courtyard during the Mid-Autumn Festival, a young Chinese woman eating spicy dumplings and Szechuan dishes at a bustling night market adorned with bright neon signs and festive decorations, a Chinese teenager eating baozi while wandering through a tranquil cherry blossom park on a crisp spring morning, or a renowned Chinese chef eating samples of an innovative fusion dish in the kitchen of a high-end, modern restaurant overlooking the city skyline.
Figure 19: Structured prompts provided to GPT-4 for generating evaluation prompts within the specific category of “Age.” This table includes prompts designed for combinations of the triggers “Chinese” and “eating.” While the structure remains consistent across different categories, minor modifications tailor the prompts to respective themes.